# EvAlignUX: Advancing UX Research through LLM-Supported Exploration of Evaluation Metrics

### v1

@@ 5.2. EvAlignUX Feature Use and Perceived Value (RQ2)
In this section, we discuss how the individual features of
EvAlignUX helped overcome different challenges in developing evaluation plans while also streamlining the thought process. In particular, three features, metrics graph, metrics list, and outcomes, are regarded as the most effective.
@@ 5.2.1. The Value of EvAlignUX Features in Enhancing UX Evaluation
We observed how the use of individual features and the combined use of different features of EvAlignUX enhances the user experience. In Figure 8, we summarized a group of prototypical questions that UX researchers asked themselves when developing their UX evaluation plans, annotated from think-aloud data during participants’ thought processes, which highlighted five perceived benefits of AI: justify, inform, support self-critique, streamline, and vision.
@@
Participants mentioned how AI could be used to Inform them knowledge in prior literature, e.g., to quickly “suggest relevant literature” (P15, 18), “summarize existing evaluation methods based on the research context” (P3, P11), supporting researchers in understanding appropriate metrics by providing “definitions” (P8), “citations” (P3), “how it can be used” (P16), and “constructs adopted” (P17). For example, P19 combined the Index and Metrics List features. P9 initially felt that EvAlignUX did not accurately capture their intention. They proposed a version control system, but EvAlignUX recommended metrics that were more likely to measure ”machine learning efficiency,” such as accuracy. However, after changing a few indexes (e.g., adding two design novelties: graph-based version control, and automated design documentation), they found EvAlignUX suggested new metrics, such as interaction experience, overall satisfaction, perceived usability, and task success, to be much more accurate and aligned with their research.
@@
Many participants recognized the value of using AI to visualize literature networks, which help them better Justify metric choices to align with research goals. This capability was particularly appreciated as it facilitates “metric selection across various venues and disciplines” (P17), “highlight connections between metrics” (P16) to determine whether two metrics are “interrelated or distinct” (P13) and support decisions depends on “popularity and relevance in the field” (P9) in a way that “aligns with their (my) research goal” (P6). For example, P9 used Metrics Graph and Outcome together. P9 designed a conversational agent for judges. After reviewing the graph, they found that user satisfaction, accuracy, and task completion were connected, which motivated him to think about the relationship between these three metrics. However, he found that while certain metrics aligned with their research goals, not all findings were applicable to their specialized context. They noted, no emotions or diversity should be included as my evaluation outcomes; otherwise, the judges will be easily biased.
@@
Participants recognized the value of using AI to help them better Self-Critique during their UX evaluation plan development process. This capability was particularly appreciated as it facilitates “prompt feedback” (P1) with “new perspectives” (P15), acting as a “teammate” (P6, P16) or “partner” (P18) in scholarly settings, to comment on their proposed plans are “effectiveness and limitations” (P13). One participant also mentioned about its potential to support ”overcoming social inhibitions, like hesitancy in seeking advice or posing questions to peers too frequently” (P16). As P4 commented that, ”It’s nice also to have EvAlignUX where you can throw ideas around, and then it will guide you directly in your research project, what to think about—not just a chat with somebody, which can sometimes be embarrassing.”
@@
Participants also recognized the value of using AI to reduce cognitive load, which helps them better Streamline their UX evaluation development to align with dynamic project timelines. This capability was particularly appreciated as it ”simplify decision-making processes in UX design” (P16) to determine whether new updates are ”directly contributing to the UX” (P13) of their original research goal or not. For example, P4 said, ”Generally if I need to propose a UX plan, I probably need to spend like two days at least to collect all the paper. I believe your system can actually help us [HCI researchers] to shorten the time. Because every paper here is coming from valid sources, so we can trust the result.” Furthermore, participants also shared the value of using the system for re-structuring their thought-process of the UX plan development. Figure 9 is an example illustrating how P2 modified the literature review process by using AI to conduct the literature review and brainstorm metrics earlier in the process.
@@
Nearly all participants recognized the value of using AI to promote early detection of risks, which helps them better Vision their UX planning. This capability was particularly appreciated as it facilitates “spotting potential risks before they turn into real headaches after the study is done” (P5), and supports decisions based on “minimizing unexpected challenges”(P9). As P10 said mentioning the Risk feature, ”It doesn’t simply show a list of risks but encourages users to click the associated link to review the actual report.”
@@ 5.2.2. Enriched Thought-Processes with the Support of EvAlignUX
Besides how participants used the EvAlignUX features differently, they also presented diverse thought processes during their development of UX evaluation plans. When they updated their thought processes, they all revised their literature review process by incorporating relevant EvAlignUX features. More than half of the participants expressed that AI facilitates their iterative loops during developing an evaluation plan (8 participants). For instance, Figure 10. illustrates P13, a UX researcher with 1-2 years experience, revised thought process after using EvAlignUX , showing AI’s role as a supportive teammate in this process. The AI-suggested metrics, as highlighted by the purple boxes, prompted participants to critically assess the relevance and appropriateness of AI’s recommendations, ensuring alignment with system design goals and expected results. This iterative reflection, where P13 evaluate the added value of AI’s input, emphasizes AI’s role in fostering critical thinking and refinement.
Similarly, Figure 11. demonstrates how P11, with 3-4 years of UX experience, views AI can benefit the process of the iterative metric selection process.
@@
We also observed that UX risks and outcomes were added into the thought process, regardless of participants’ experience level. Overall, 42% of participants incorporated new components into their UX evaluation development thought processes after using EvAlignUX . Notably, these included ”Propose measurements” (3 junior UX-level participants, mixed research stages), ”UX risks” (mixed UX levels, mixed research stages), and ”Research outcomes” (mixed UX levels, mixed research stages). These additions reflect a newly gained awareness of potential challenges and impacts in their UX evaluation plan development. For example, Figure 12 shows how P7, a UX researcher with 1-2 years of experience, added two new considerations: defining the ideal outcome of the evaluation and identifying potential risks associated with the chosen measurements to their thought process after using EvAlignUX .
@@ 5.2.3. The Significant Use of Hypothesized Outcome and Metrics Exploration Features
Based on these observations, we further analyzed the relationship between users’ system log variables (e.g., number of times indexes were regenerated, number of metrics selected via the Metrics List, number of metrics selected or deselected via Metric Graphs, number of outcomes clicked, number of risks clicked) and their perceptions of five different features across 11 perception variables (e.g., enjoyment, intention, usefulness, cognitive load, inspiration), see Section 4.3.3 for measurements. A total of 55 pairs were examined, with 6 showing significant results. All significant results are presented in Table 3.
@@
The result showed that selecting more potential outcome items (i.e., previous research findings) is associated with higher user ratings of the Outcome feature, both enjoyment (r = 0.57, 95% CI [0.28, 0.82], p ¡ 0.01∗∗) and in future use (i.e., intention to use: r = 0.40, 95% CI [0.05, 0.66], p ¡ 0.05∗). Additionally, users rated this feature more positively in terms of its contribution to revising their project idea (r = 0.45, 95% CI [0.05, 0.76], p ¡ 0.05∗), UX evaluation plan (r = 0.44, 95% CI [0.08, 0.71], p ¡ 0.05∗), and UX outcomes (r = 0.49, 95% CI [0.16, 0.75], p ¡ 0.05∗). The findings from relevant HCI domains encouraged them to think about questions such as what demographic is left in the original plan? (P5) What factor was not considered yet? (P14) These thoughts made them to refer back to their original project idea, UX evaluation plan, and UX outcomes. P4 selected 15 potential outcomes that were derived from prior work, they commented that for instance, ”measuring trust and confidence in the chatbot’s abilities and context understanding” was ”exactly what they were looking for and this point really hit the point”. P4 further commented, they liked this feature because they felt ”the generation understands my initial thoughts about the project… It helps me finalize the details of my plan.”
@@ 5.3. EvAlignUX Perception Varying by Research Background(RQ3)
In this section, we examined the relationship between users’ backgrounds and their perceptions of 5 different features across 9 perception variables (e.g., enjoyment, intention, usefulness, cognitive load, inspiration, see measurements in section 4.3.3), aiming to identify the users who benefit the most from our system and the users we still need to improve support, with 7 showing significant results. All significant results are shown in Table 4.
@@ 5.3.1. The Extent of Perceived Inspiration Varies by Users’ Research Experience
Overall, the perceived value of providing inspiration through EvAlignUX was significantly negatively correlated with participants’ background, with the exception of the Outcome feature. Namely, participants with less UX experience provided higher scores for perceived inspiration, except with the Outcome feature. Despite this, all participants rated the inspirational measure highly (M = 3.58–4.00, SD = 0.94–1.33).
@@
For Index, some participants with less experience shared that, the feature helps to structure research ideas using relevant terminologies as sometimes they ”just have a vague thought without knowing the exact terms” (P9). P6, with 1 year of UX experience, shared that the feature is a ”Very good at pinpointing important aspects, like user-centered design and stakeholders (researchers). It did a good job of checking what the user intended. I wouldn’t say anything was wrong or needs changing. Everything was spot on. There are certain things you can add, but they’re already in line with what was checked before.” For experienced researchers like P2, she mentioned that the feature helps to distinguish ”umbrella terms” when they only have a vague understanding of what they are measured: ”It lowered the effort to identify the literature that is actually related. For instance, I saw three terms, knowledge acquisition, knowledge retention, and knowledgability. If I just use the wrong term, I may think that no one has done it before, but it turns out people have already done it.” While experienced users acknowledged its utility for combining suggested indexes in their paper searches, they felt it lacked novelty, as many of the indexes were already familiar (P2), ”I am looking for something that I have no knowledge about.” Instead, they tend to ”self-validation whether they are on the right track” (P4) instead of seeking inspiration for relevant knowledge.
@@
For Metrics List and Metrics Graph, less experienced users valued it for revealing ”metrics [they] hadn’t considered” (P7). As P3 (1-2 years of UX experience) mentioned, ”Because generally if I need to do that (on my own), I probably need to spend like 2 days at least to collect all the paper, which is really bad. I think EvAlignUX can actually help us to shorten the time. Because every paper here is coming from valid sources, such as top conferences, so we can trust the result.” Similarly, P6 said, the graph gives her serendipity, ”An aerial view of how these evaluation schemes tie together, like from usefulness to ease of use, helps when writing the paper or an evaluation scheme. Certain things in usefulness and trust are often correlated, as shown by the thickness of the line, which is inspiring.” The graphs help researchers to identify is their chosen metrics are “novel”, because the size of the nodes represent the frequency a metric is used in a research domain; this also help them to discover some “under-studied” metric in a domain; also for commonly used metric, they got to know what are the common practices. As P7 said, ”I feel like changing my selections based on the graph resulted in better and more related contributions.”
@@
This pattern extended to the Risk feature, where P6 with 1-2 UX research experience found it interesting to identify ”risks that directly related to the system.” More experienced researchers, like P1, mentioned that ”The risks mentioned here are not hardlines, meaning that if users are not followed, then the system will fail. For instance, if the evaluation plan doesn’t get the IRB approval, they cannot continue to the next step; whereas the potential risk information here is more likely to be soft suggestions, which may not contribute to actual revision.”
@@
Interestingly, inspiration from the Outcome feature did not vary with UX experience, aligning with participants drawing of thought-process, where outcomes were often missing from initial UX plans regardless of participants’ UX experience. The overall high mean score of rated inspiration from Outcome suggests the feature resonated across all users, though for different reasons based on experience level. Several participants with less UX experience found the Outcome feature useful for refining their research plans, noting how it helped them ”rephrase [their] outcome to make it more detailed” (P13). P7 said, “I haven’t thought about it before, but now that I saw it here, that can be an useful contribution…It’s really cool that they come from actual papers and aren’t fabricated or made up.” Whereas, P4, with 3-4 years of UX experience, noted that “without going into too much like spending a day thinking about it (UX outcomes). I think that it (your suggested UX outcomes) is a really great restatement of my original one.”

### v2

@@ 5.2. EvAlignUX Feature Use and Perceived Value (RQ2)
In this section, we discuss how features of EvAlignUX helped overcome various challenges in developing evaluation plans while also streamlining the thought process. In particular, three features, metrics graph, metrics list, and outcomes, were regarded as the most effective.
The image presents a UX Evaluation Development Questions Bank, divided into two main sections: General Questions and Questions Related to AI Support. These sections are further categorized into key areas of UX development: Research Goal, Literature Review, Evaluation Metrics, Research Impacts, External Feedback, and System Prototype. The General Questions section provides foundational queries for traditional UX evaluation, while the Questions Related to AI Support section includes prompts considering AI as the research assistant. The questions cover various stages of UX research and design, guiding users through research planning, evaluation, and the integration of AI-generated insights.
@@ 5.2.1. The Value of EvAlignUX Features in Enhancing UX Evaluation
We observed how the use of individual features and the combined use of different features in EvAlignUX enhanced the user experience. In Figure 9, we summarized a set of prototypical questions that UX researchers asked themselves while developing their UX evaluation plans, annotated from think-aloud data collected during participants’ thought processes, which highlighted five perceived benefits of AI: justify, inform, support self-critique, streamline, and vision.
@@
Participants mentioned that AI could be used to inform them by providing knowledge from prior literature, e.g., to quickly “suggesting relevant literature” (P15, 18), “summarizing existing evaluation methods based on the research context” (P3, P11). AI also supported researchers in understanding appropriate metrics by offering “definitions” (P8), “citations” (P3), “how it can be used” (P16), and information on “constructs adopted” (P17). For example, P19 combined the Index and Metrics List features. P9 initially felt that EvAlignUX did not accurately capture their intentions. They proposed a version control system, but EvAlignUX recommended metrics that were more likely to measure ”machine learning efficiency,” such as accuracy. However, after modifying a few indexes (e.g., adding two design novelties: graph-based version control, and automated design documentation), they found that EvAlignUX suggested new metrics, such as interaction experience, overall satisfaction, perceived usability, and task success, that were much more accurate and aligned with their research.
@@
Many participants recognized the value of using AI to visualize literature networks, which help them better justify metric choices to align with research goals. This capability was particularly appreciated because it facilitates “metric selection across various venues and disciplines” (P17), “highlight connections between metrics” (P16) to determine whether two metrics are “interrelated or distinct” (P13) and supports decisions based on “popularity and relevance in the field” (P9) in a way that “aligns with their (my) research goal” (P6). For example, P9 used Metrics Graph and Outcome together. P9 designed a conversational agent for judges. After reviewing the graph, they found that user satisfaction, accuracy, and task completion were connected, which motivated them to think about the relationship between these three metrics. However, they found that while certain metrics aligned with their research goals, not all findings were applicable to their specialized context. They remarked that, no emotions or diversity should be included as my evaluation outcomes; otherwise, the judges would be easily biased.
@@
Participants recognized the value of using AI to help them better self-critique during the development of their UX evaluation plans. This capability was particularly appreciated as it facilitates “prompt feedback” (P1) and provides “new perspectives” (P15), acting as a “teammate” (P6, P16) or “partner” (P18) in scholarly settings by commenting on the “effectiveness and limitations” of their proposed plans (P13). One participant also mentioned its potential to support ”overcoming social inhibitions, such as hesitancy in seeking advice or posing questions to peers too frequently” (P16). As P4 commented that, ”It’s nice also to have EvAlignUX where you can throw ideas around, and then it will guide you directly in your research project, what to think about—not just a chat with somebody, which can sometimes be embarrassing.”
@@
Participants also recognized the value of using AI to reduce cognitive load, which helps them better streamline their UX evaluation development and align it with dynamic project timelines. This capability was particularly appreciated as it ”simplify decision-making processes in UX design” (P16) to determine whether new updates are ”directly contributing to the UX” (P13) of their original research goal or not. For example, P4 said, ”Generally if I need to propose a UX plan, I probably need to spend like two days at least to collect all the papers. I believe your system can actually help us [HCI researchers] to shorten the time. Because every paper here comes from valid sources, so we can trust the results.” Furthermore, participants also shared the value of using the system for re-structuring their thought process of the UX plan development. Figure 10 is an example illustrating how P2 modified the literature review process by using AI to conduct the literature review and brainstorm metrics earlier in the process.
This graph compares P2’s thought process using an original sequential workflow versus a more efficient, AI-assisted parallel workflow. In the original sequential flow, tasks are completed one after the other. It starts with identifying research questions (RQs), followed by testing the feasibility of implementation, then moving on to a design study, and continues with additional iterative steps. In the new parallel workflow, AI reduces manual effort, allowing certain tasks to occur simultaneously. Specifically, AI is used to conduct literature reviews and brainstorm metrics, which enables the design study and the feasibility test to happen in parallel. This approach, by lowering the effort required, facilitates faster updates to evaluation plans.
@@
Nearly all participants recognized the value of using AI to promote early detection of risks, which helps them better envision their UX planning. This capability was particularly appreciated as it facilitates “spotting potential risks before they turn into real headaches after the study is done” (P5), and supports decisions based on “minimizing unexpected challenges”(P9). As P10 said mentioning the Risk feature, ”It doesn’t simply show a list of risks but encourages users to click the associated link to review the actual report.”
The figure shows the differences in P13’s approach to developing UX evaluation plans without and with EvalignUX, focusing on the “Index” and “Metric Graph” features. On the upper left side, a simple flowchart illustrates P13’s initial thought process, which begins with deciding whether to use a baseline for comparison and then moves through selecting different study conditions and measures—whether qualitative, quantitative, or mixed. After integrating AI as a “teammate,” the process expands to include steps like reading the AI’s references, examining the AI’s suggested metrics, and determining their relevance to the system’s design goals. On the right, a snapshot of the Metric Graph interface shows how P13 repeatedly refined the indexes by using AI’s suggestions to uncover novel insights.
@@ 5.2.2. Enriched Thought-Processes with the Support of EvAlignUX
Besides how participants used the EvAlignUX features differently, they also demonstrated diverse thought processes during their development of UX evaluation plans. When updating their thought processes, they all revised their literature review process by incorporating relevant EvAlignUX features. More than half of the participants expressed that AI facilitates iterative loops during the development of an evaluation plan (8 participants). For instance, Figure 11 illustrates P13, a UX researcher with 1-2 years experience, revised thought process after using EvAlignUX , showing AI’s role as a supportive teammate in this process. The AI-suggested metrics, as highlighted by the purple boxes, prompted participants to critically assess the relevance and appropriateness of AI’s recommendations, ensuring alignment with system design goals and expected results. This iterative reflection, where P13 evaluates the added value of AI’s input, emphasizes AI’s role in fostering critical thinking and refinement. Similarly, Figure 12. demonstrates how P11, with 3-4 years of UX experience, views AI as beneficial to the iterative metric selection process.
The image depicts P11’s evolving thought process while utilizing AI as a teammate to refine UX evaluation plans, focusing on the “Metrics List” feature. The diagram illustrates how P11 initially considered comparing techniques, study types, and study factors before performing tasks and measuring outcomes. The process then branches into different types of measurements, including objective and subjective measurements, while also considering ethical aspects. The blue lines indicate where AI’s involvement is perceived to add benefits during the iterative process of metric selection. AI aids P11 by providing references on how certain metrics, like “engagement”, can be measured using surveys and system logs. Additionally, P11 incorporated AI-suggested risks, noting that AI helped refine metrics by highlighting potential risks and emphasizing the importance of collecting objective measurements to address limitations.
The graph illustrates P7’s evolving thought process for developing a UX evaluation strategy with EvalignUX. Initially, P7 considers broad questions, such as the key research questions and which user interactions or features to observe—before gradually narrowing down to a specific evaluation goal. Along the way, P7 refines the plan using the Outcome and Risk features. By selecting checkboxes from “Results from Prior Studies” and “Potential Risks,” P7 reviews previous findings for inspiration, extracts meaningful insights, and acknowledges potential downsides. The revised UX outcomes now include both improvements, like enhanced conversation quality, better engagement, and ease of use, and potential pitfalls, such as user disengagement or misinterpretation of AI guidance. Arrows and speech-bubble-like shapes connect these steps, showing the progression from defining evaluation goals and desired user behaviors to selecting valuable prior results and understanding associated risks. This balanced approach demonstrates that P7 is considering both positive and negative outcomes to shape research values and identify areas for improvement.
@@
We also observed that UX risks and outcomes were incorporated into participants’ thought processes, regardless of their experience level. Overall, 42% of participants incorporated new components into their UX evaluation development thought processes after using EvAlignUX . Notably, these included ”Propose measurements” (3 junior UX-level participants, mixed research stages), ”UX risks” (mixed UX levels, mixed research stages), and ”Research outcomes” (mixed UX levels, mixed research stages). These additions reflect a newly gained awareness of potential challenges and impacts in their UX evaluation plan development. For example, Figure 13 shows how P7, a UX researcher with 1-2 years of experience, added two new considerations: defining the ideal outcome of the evaluation and identifying potential risks associated with the chosen measurements to their thought process after using EvAlignUX .
@@ 5.2.3. The Significant Use of Hypothesized Outcome and Metrics Exploration Features
Based on these observations, we further analyzed the relationship between users’ system log variables (e.g., the number of times indexes were regenerated, the number of metrics selected via the Metrics List, the number of metrics selected or deselected via Metric Graphs, the number of outcomes clicked, and the number of risks clicked) and their perceptions of five different features across 11 perception variables (e.g., enjoyment, intention, usefulness, cognitive load, inspiration), see Section 4.3.3 for measurements. A total of 55 pairs were examined, with 6 showing significant results. All significant results are presented in Table 7.
@@
Table 3 summarizes how participants’ feature uses based on system logs (e.g., the number of metrics they added using a metrics graph or the number of results they selected) relate to their perceptions of their perceived values of the features. The dependent variables like “Cognitive Load,” “Enjoyment,” “Intention to Use,” “Contribution to UX Evaluation,” and “Contribution to UX Outcomes,” each with a mean, standard deviation (SD), correlation coefficient (r), and 95% confidence interval. Significant p-values are indicated, with **p ¡ 0.01 and *p ¡ 0.05, showing that greater engagement with the Outcome feature (i.e., selecting results) is associated with higher perceived enjoyment, greater intention to use, and a stronger sense that the features contribute positively to UX evaluation and outcomes. For the Graph feature, the correlation between the number of metrics added by graph and perceived cognitive load is negative, suggesting that the Graph feature usage may be linked to a reduction in perceived mental effort.
The results showed that selecting more potential outcome items (i.e., previous research findings) is associated with higher user ratings of the Outcome feature, both enjoyment (r = 0.57, 95% CI [0.28, 0.82], p ¡ 0.01∗∗) and intention to use (r = 0.40, 95% CI [0.05, 0.66], p ¡ 0.05∗). Additionally, users rated this feature more positively in terms of its contribution to revising their project idea (r = 0.45, 95% CI [0.05, 0.76], p ¡ 0.05∗), UX evaluation plan (r = 0.44, 95% CI [0.08, 0.71], p ¡ 0.05∗), and UX outcomes (r = 0.49, 95% CI [0.16, 0.75], p ¡ 0.05∗). The findings from relevant HCI domains encouraged participants to consider questions such as what demographic is left in the original plan? (P5) What factor was not considered yet? (P14) These thoughts prompted them to refer back to their original project idea, UX evaluation plan, and UX outcomes. P4 selected 15 potential outcomes that were derived from prior work, they commented that for instance, ”measuring trust and confidence in the chatbot’s abilities and context understanding” was ”exactly what they were looking for and this point really hit the point”. P4 further commented, they liked this feature because they felt ”the generation understands my initial thoughts about the project… It helps me finalize the details of my plan.”
@@ 5.3. EvAlignUX Perception Varying by Research Background (RQ3)
In this section, we examined the relationship between users’ backgrounds and their perceptions of five different features across nine perception variables (e.g., enjoyment, intention, usefulness, cognitive load, inspiration, see measurements in Section 4.3.3), aiming to identify which users benefit the most from our system and which users require additional support. Seven significant results were found. All significant results are shown in Table 4.
@@
Table 4 shows how participants’ backgrounds, measured by their self-reported UX experience, relate to their perceived inspiration drawn from system features. The table includes features, Index, Metrics list, Metrics Graph, Outcome, and Risk, along with their corresponding perceived inspiration scores (mean and standard deviation), correlation coefficients (r), and 95% confidence intervals. Statistically significant results are denoted by * (p ¡ 0.05) and ** (p ¡ 0.01). For the Index, Metrics list, Metrics Graph and Risk features, there are significant negative correlations, suggesting that participants with more UX experience rated these features as less inspiring, whereas those with less experience found them more inspiring. For the Outcome feature, there is no significant difference between participants with junior UX experience and senior UX experience. In summary, this table illustrates how users’ prior UX experience levels can influence how inspirational they find certain system features.
@@ 5.3.1. The Extent of Perceived Inspiration Varies by Users’ Research Experience
Overall, the perceived value of providing inspiration through EvAlignUX was significantly negatively correlated with participants’ background, with the exception of the Outcome feature. Specifically, participants with less UX experience provided higher scores for perceived inspiration, except with the Outcome feature. Despite this, all participants rated the inspirational measure highly (M = 3.58–4.00, SD = 0.94–1.33).
@@
For Index, some participants with less experience shared that the feature helps to structure research ideas using relevant terminologies as sometimes they ”just have a vague thought without knowing the exact terms” (P9). P6, with 1 year of UX experience, shared that the feature is a ”very good at pinpointing important aspects, like user-centered design and stakeholders (researchers). It did a good job of checking what the user intended. I wouldn’t say anything was wrong or needs changing. Everything was spot on. There are certain things you can add, but they’re already in line with what was checked before.” For experienced researchers like P2, she mentioned that the feature helps to distinguish ”umbrella terms” when they only have a vague understanding of what is being measured: ”It lowered the effort to identify the literature that is actually related. For instance, I saw three terms, knowledge acquisition, knowledge retention, and knowledgability. If I just use the wrong term, I may think that no one has done it before, but it turns out people have already done it.” While experienced users acknowledged its utility for combining suggested indexes in their paper searches, they felt it lacked novelty, as many of the indexes were already familiar (P2), ”I am looking for something that I have no knowledge about.” Instead, they tend to ”self-validate whether they are on the right track” (P4) rather than seek inspiration for relevant knowledge.
@@
For Metrics List and Metrics Graph, less experienced users valued these features for revealing ”metrics [they] hadn’t considered” (P7). As P3 (with 1-2 years of UX experience) mentioned, ”Because generally if I need to do that (on my own), I probably need to spend like 2 days at least to collect all the paper, which is really bad. I think EvAlignUX can actually help us to shorten the time. Because every paper here is coming from valid sources, such as top conferences, so we can trust the result.” Similarly, P6 said that the graph gives her serendipity, ”An aerial view of how these evaluation schemes tie together, like from usefulness to ease of use, helps when writing the paper or designing an evaluation scheme. Certain things in usefulness and trust are often correlated, as shown by the thickness of the line, which is inspiring.” Participants shared that the graph helped them explore metrics that were “novel” (P3), as they noticed smaller nodes, disconnected nodes, or nodes with unfamiliar names. The graph made them feel more interested in discovering “under-studied” (P6) metrics in a domain; and for commonly used metrics, they felt they gained insights into common practices (P7, P12). As P7 shared, ”I feel like changing my selections based on the graph resulted in better and more related contributions.”
@@
This pattern extended to the Risk feature. P6, with 1-2 years of UX research experience, found it interesting that the feature helped identify ”risks that directly related to the system.” More experienced researchers, like P1, mentioned that ”The risks mentioned here are not hardl ines, meaning that if users are not followed, then the system will fail. For instance, if the evaluation plan doesn’t receive IRB approval, they cannot continue to the next step; whereas the potential risk information here is more likely to be soft suggestions that may not contribute to actual revision.”
@@
Interestingly, inspiration from the Outcome feature did not vary with UX experience, aligning with participants drawing of thought-process, where outcomes were often missing from initial UX plans regardless of participants’ UX experience. The overall high mean score for rated inspiration from Outcome suggests that the feature resonated across all users, though for different reasons based on experience level. Several participants with less UX experience found the Outcome feature useful for refining their research plans, noting how it helped them ”rephrase [their] outcome to make it more detailed” (P13). P7 said, “I haven’t thought about it before, but now that I saw it here, that can be an useful contribution… It’s really cool that they come from actual papers and aren’t fabricated or made up.” In contrast, P4, with 3-4 years of UX experience, noted that “without going into too much like spending a day thinking about it (UX outcomes). I think that it (your suggested UX outcomes) is a really great restatement of my original one.”